# Hybrid Deep Learning Model to Estimate Cognitive Effort from fNIRS Signals in Educational Game Playing

### v2

@@ 6. Discussion
For RQ1, the results suggest that the DeepNet model performed better than traditional machine learning models. It also shows that using CNN, LSTM, and BiLSTM alone does not give the best results for score prediction using fNIRS. The hybrid CNN-GRU model achieved the highest accuracy (73.08%). This suggests that a combination of CNN for feature extraction and GRU helps improve classification performance.
The inclusion of fNIRS signals with DeepNet enabled the estimation of cognitive metrics such as RNE and RNI from predicted performance scores. For RQ2 these derived measures help interpret learners’ cognitive effort, involvement, and efficiency during educational gameplay.
These findings highlight the importance of personalized learning design. We believe the behavior of RNE and RNI is influenced by several cognitive and physiological factors. For example, CE dropped down after a post-break, which aligns with the hypothesis that rest helps restore executive functioning. Also, as participants became more familiar with the task structure, learning progression probably helped to improve efficiency and involvement. These transitions suggest that both task familiarity and rest contribute to variations in neural efficiency and involvement over time.
For RQ3 these derived measures help interpret learners’ cognitive effort, involvement, and efficiency during educational gameplay.By visualizing predicted and actual scores alongside cognitive effort, instructors and researchers can adjust instructional strategies and interventions more precisely. Results suggest that RNE and RNI are not static but respond to breaks, fatigue, and learning progression. Breaks consistently led to performance recovery, as seen in increased RNE post-break. For RQ3 these derived measures help interpret learners’ cognitive effort, involvement, and efficiency during educational gameplay. Even with moderate accuracy, the predicted CE trends followed actual RNE/RNI patterns. Since we use aggregated and z-scored values for CE estimation, occasional false positives/negatives have a limited impact on the overall cognitive state trends. The derived RNE/RNI values smooth out individual prediction errors. Also, to observe cognitive effort patterns, we used a Cartesian plot. We observed the predicted values most of the time aligned with the actual values in the same quadrant of the Cartesian plot. This spatial overlap indicates that the model was able to capture similar patterns of cognitive effort behavior. Therefore, the proposed model not only supports cognitive effort estimation but also aids in the interpretation of cognitive state trends across different learning segments.
However, future improvements can reduce this noise further. Our goal was not to maximize classification accuracy alone, but to enable interpretable CE estimation using RNE and RNI. The CNN-GRU model offered a good tradeoff between accuracy and signal pattern learning. We plan to explore more robust or adaptive architectures in future work. CE is highly individualized. While our model captures general trends, it does not yet account for individual baselines or variability in fNIRS responses. Adaptive or personalized calibration may improve robustness in future work.
@@
Real-time implementation was beyond the scope of this study, but the CNN-GRU model we used is lightweight. So, it should be able to make predictions in real-time. In the future, deploying this system in real-time during learning sessions could help detect fatigue and adjust task difficulty dynamically, contributing to more effective and personalized educational experiences. To improve generalizability, future work will explore data augmentation (e.g., jittering, time warping of signals) and transfer learning from larger fNIRS datasets. We also plan to extend the study to diverse populations in classroom environments to validate the model. Additionally, adjusting the duration of learning tasks, such as extending or shortening session lengths, may provide further insight into how cognitive effort and motivation evolve. Exploring the impact of different types of learning materials (e.g., visual vs. interactive, gamified vs. traditional) on RNE and RNI could also uncover new instructional design principles.
@@ 7. Conclusion
This study demonstrates a method for estimating and interpreting cognitive effort through performance scores predicted by DeepNet and oxygenated hemoglobin ($\Delta HbO$) in the brain collected using fNIRS. In this work, we proposed a hybrid model to predict performance scores in an educational game. We collected $\Delta HbO$ using fNIRS and fed it into our CNN-GRU model for the prediction. This predicted score was used to estimate cognitive effort by measuring RNE and RNI. Our hybrid model showed 73.08% accuracy with 82% F1-Score. After estimating RNE and RNI, we observed that the predicted RNE and RNI closely follow the actual RNE and RNI. This approach paves the way for a personalized educational setup that can dynamically adapt to learners’ cognitive states, potentially transforming classroom and remote educational settings. By deriving RNE and RNI from predicted performance and effort, we reveal meaningful trends tied to cognitive states. The findings highlight the model’s potential in adaptive learning contexts and highlight future directions in real-time neuro-adaptive systems.
Acknowledgments
We wish to express our gratitude to the study participants and lab members.

### v3

@@ 6. Discussion
For $RQ_{1}$, the results suggest that the DeepNet model performed better than traditional machine learning models.The hybrid CNN-GRU model achieved the highest accuracy (73.08%). This suggests that a combination of CNN for feature extraction and GRU helps improve classification performance.
For $RQ_{2}$, we conducted a three-stage interpretability analysis combining PCA, GRU feature correlation, and SHAP values. We first mapped PCA components to anatomical regions and found that PC2 was dominated by signals from the LPFC. We then observed that GRU Feature 6, which had the strongest correlation with PC2, likely encoded LPFC-related signals. SHAP analysis further identified GRU Features 0 and 3 as most influential in final decision-making, both of which were also correlated with PC2. These findings suggest that LPFC activity contributed significantly to performance prediction and, by extension, to our estimation of cognitive effort.
For $RQ_{3}$, the segment-wise results show that RNE and RNI values calculated from predicted scores closely match the actual cognitive effort patterns observed in the fNIRS signals. While there were some variations in MAE across participants, particularly for P8, the Pearson correlation remained high ($r>0.94$), indicating that the predicted and actual trends were well aligned. This suggests that the model may not always predict the exact value but still captures the overall direction of change.
The scatter plots in Figure 5 support this finding. Most points are close to the diagonal, which indicates that the predicted values follow both the order and the scale of actual effort levels. Participant P16 showed nearly perfect alignment, suggesting that the model can perform reliably even at the individual level.
Although the performance prediction accuracy was moderate, around 73%, the calculated RNE and RNI still matched the actual values. This is because RNE and RNI depend on the relationship between performance and brain signals, not just the predicted score label. In our method, we used standardized performance and neural activity. As a result, even when the predicted score was not exact, the combined trend remained consistent. This 73% accuracy is influenced by several factors such as the inherent noise and variability of fNIRS signals, the relatively small participant sample, and the subtle nature of hemodynamic responses during low-stakes educational tasks. Despite this, the RNE and RNI metrics remain robust because they rely on trends between brain activation and performance rather than exact score classification.
This conclusion is further supported by the low MAE and high correlation. These results show that if performance can be predicted in real time and fNIRS data are available, it is possible to estimate cognitive effort for a given time window. This remains achievable even when classification accuracy is moderate, as long as the predicted scores reflect the interaction between performance and neural activation.
@@
Real-time implementation was beyond this study’s scope, but our lightweight CNN-GRU model is suitable for real-time deployment. Cognitive effort was estimated post-hoc using predicted scores and fNIRS signals. In future work, we plan to build a real-time system that estimates CE during learning using live fNIRS data and dynamic model outputs. To improve robustness, we aim to incorporate personalized calibration and enhance generalizability through data augmentation and transfer learning. We also plan to explore how session duration and content type influence CE and extend validation to broader populations and real-world classrooms to support adaptive learning systems grounded in neural feedback.
@@ 7. Conclusion
This study demonstrates a method for estimating and interpreting cognitive effort through performance scores predicted by DeepNet and oxygenated hemoglobin ($\Delta HbO$) in the brain collected using fNIRS. In this work, we proposed a hybrid model to predict performance scores in an educational game. We collected $\Delta HbO$ using fNIRS and fed it into our CNN-GRU model for the prediction. This predicted score was used to estimate cognitive effort by measuring RNE and RNI. Our hybrid model showed 73% accuracy with 82% F1-Score. After estimating RNE and RNI, we observed that the predicted RNE and RNI closely follow the actual RNE and RNI. This approach paves the way for a personalized educational setup that can dynamically adapt to learners’ cognitive states, potentially transforming classroom and remote educational settings. By deriving RNE and RNI from predicted performance and effort, we reveal meaningful trends tied to cognitive states. The findings highlight the model’s potential in adaptive learning contexts and highlight future directions in real-time neuro-adaptive systems.
Acknowledgment
We express our gratitude to the study participants and lab members. We also thank the National Science Foundation for its support (#222266122226632222661-2222663, #$2321274$, and #$2426003$).